OpenStudy (anonymous):

Find the measure of Angle 8. measure of angle 7 = 2x+15 measure of angle 8 = 3x

OpenStudy (agent47):

im confuzzled o.O

OpenStudy (agent47):

there was a graph attached, wasn't there..

OpenStudy (anonymous):

yup, but not helpful at alll. lol

OpenStudy (agent47):

yes it is...2*angle7=2(2x+15) 2*angle8=2*3x 2(2x+15)+6x=360

OpenStudy (agent47):

4x+30+6x=360 10x=360-30 10x=330 x=33

OpenStudy (agent47):

angle8=3*33=99

OpenStudy (anonymous):

thanks!